OpenStudy (anonymous):

Please help!!!! How do you verify cot(x-pi/2) = -tan(x)

mathslover (mathslover):

\(\cot \theta = \cfrac{\cos \theta}{\sin \theta}\) Therefore, \(\cot \left( x- \pi /2 \right) = \cfrac{\cos (x - \pi/2)}{\sin (x- \pi /2)}\)

mathslover (mathslover):

Now, what is the formula for : \(\cos (x- \pi/2) \) and \(\sin (x-\pi/2)\) ?

OpenStudy (anonymous):

Uhm... Shoot I can't remember what the formula for it is! Give me a second

OpenStudy (anonymous):

Sin(π/2-x) = cosx cos(π/2-x) = sinx?

mathslover (mathslover):

Yep. that's right, but notice that we have : \(\cos (x - \pi/2) \) and \(\sin (x -\pi /2)\) instead of \(\cos(\pi/2 - x)\) and \(\sin (\pi/2 -x)\) .

mathslover (mathslover):

So, let us play with what we have. \(\cos(x - \pi/2)\) can be written as : \(\cos (-(\pi/2 - x)) \) , right?

OpenStudy (anonymous):

I think so

mathslover (mathslover):

Okay. So, remember this that : \(\cos (-\theta) = \cos (\theta)\) while \(\sin (-\theta) = - \sin (\theta)\)

OpenStudy (anonymous):

So, that would make it cos(x)/-sin(x) right?

mathslover (mathslover):

Therefore, I can write : \(\cos(-(x - \pi/2) ) \) as \(\cos (\pi/2 -x)\) and \(\sin (-(x-\pi/2)) \) as \( - \sin (\pi/2 - x) \)

mathslover (mathslover):

What we have now is : \(\cfrac{\cos (\pi/2 - x)}{- \sin (\pi/2 -x) }\)

mathslover (mathslover):

Use this now : Sin(π/2-x) = cosx cos(π/2-x) = sinx what do u get?

OpenStudy (anonymous):

sin(x)/-cos(x) ?

mathslover (mathslover):

Excellent ! So, you get : \(\cfrac{\sin (x) }{-\cos (x)}\) What is \(\cfrac{\sin (x)}{\cos (x)}\) ?

OpenStudy (anonymous):

tan(x)

OpenStudy (anonymous):

So, that would mean that it would be the -tan(x), right?

mathslover (mathslover):

good, so there was a minus in denominator, it becomes tan(x) GOOD WORK!

OpenStudy (anonymous):

So, all together it would be cos(x-π/2)/sin(x-π/2) cos(π/2-x)/-sin(π/2-x) sin(x)/-cos(x) = -tan(x) -tan(x)=-tan(x) Would this be the work for it then?

mathslover (mathslover):

Yeah, right!

OpenStudy (anonymous):

Thank you so much for your help!!!

mathslover (mathslover):

You're welcome. Also, Welcome to OpenStudy!

OpenStudy (anonymous):

Thank you!! You are incredibly helpful!

mathslover (mathslover):

:-) Good to hear that I was of some help.